OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

The problem is: \[\sqrt{6}(\sqrt{3}+5\sqrt{2})\]

OpenStudy (anonymous):

Answers: A. \[3\sqrt{3}+15\sqrt{2}\] B. \[3\sqrt{2}+5\sqrt{12}\] C. \[3\sqrt{2}+10\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

\[\sqrt{6}(\sqrt{3}+5\sqrt{2})\] \[\sqrt{6}(\sqrt{3})+\sqrt{6}(5\sqrt{2})\] \[\sqrt{6*3}+5\sqrt{6*2}\] \[\sqrt{18}+5\sqrt{12}\] \[\sqrt{9*2}+5\sqrt{4*3}\] \[\sqrt{9}*\sqrt{2}+5\sqrt{4}*\sqrt{3}\] \[3\sqrt{2}+5*2*\sqrt{3}\] \[3\sqrt{2}+10\sqrt{3}\] So the answer is choice D

OpenStudy (lilg132):

the answer is C not d :p

jimthompson5910 (jim_thompson5910):

oh right, thought the last choice was D for some reason lol

jimthompson5910 (jim_thompson5910):

nice catch lilg132

OpenStudy (lilg132):

thanks :p